A) solve the following initial value problem: dy/dx= cos^2 y/4x-3 ; y(1)=π/4.


B) let F(x,y)=y cos(x^2 y^2)+y, then


(i) find the first partial derivatives Fx and Fy.

(ii) using b(i) above, find dy/dx.

(iii) if F(x,y)=0,then find dy/dx using implicit differentiation to confirm your answer in part (b) (ii) above.


1
Expert's answer
2021-08-23T16:15:11-0400

A)

"\dfrac{dy}{dx}=\dfrac{\cos^2y}{4x-3}"

"\dfrac{dy}{\cos^2y}=\dfrac{dx}{4x-3}"

"\int\dfrac{dy}{\cos^2y}=\int\dfrac{dx}{4x-3}"

"\tan y=\ln|4x-3|+C"

"y(1)=π/4, \tan \dfrac{\pi}{4}=\ln|4(1)-3|+C=>C=1"

"\tan y=\ln|4x-3|+1"

B)


"F(x,y)=y \cos(x^2 y^2)+y"

(i)

"F_x=-2xy^3\sin(x^2y^2)"

"F_y=\cos(x^2 y^2)-2x^2y^2\sin(x^2y^2)+1"

(ii)


"\dfrac{dy}{dx}=-\dfrac{F_x}{F_y}=-\dfrac{-2xy^3\sin(x^2y^2)}{\cos(x^2 y^2)-2x^2y^2\sin(x^2y^2)+1}"

"=\dfrac{2xy^3\sin(x^2y^2)}{\cos(x^2 y^2)-2x^2y^2\sin(x^2y^2)+1}"

(iii)


"F(x,y)=0=>y \cos(x^2 y^2)+y=0"

Differentiate both sides with respect to "x" and use the Chain Rule

"\dfrac{dy}{dx}\cdot \cos(x^2 y^2)-y\sin(x^2y^2)(2xy^2+2x^2y\cdot\dfrac{dy}{dx})+\dfrac{dy}{dx}=0"

Solve for "\dfrac{dy}{dx}"


"\dfrac{dy}{dx}=\dfrac{2xy^3\sin(x^2y^2)}{\cos(x^2 y^2)-2x^2y^2\sin(x^2y^2)+1}"

The answers are the same.